The graph of a quadratic function has x-intercepts -10 and 2. Write a quadratic equation in standard form that has these roots.

Thank you

Answer by ikleyn(52778)

.
f(x) = a*(x-(-10))*(x-2) = a*(x+10)*(x-2) = a*(x^2 +8x - 20) = ax^2 + 8ax -20a,      (*)

where a is an arbitrary real number.

It is the most general standard form of such a function.

If you want to have an equation, equate (*) to zero:

ax^2 + 8ax - 20a = 0.